OpenStudy (toxicsugar22):

Caffenie leaces the body at a rate of 15.6% each hou. your first cup of coffee in the morning has 100 mg of caffenie How long will it take before you have 50 mg of that caffenie in your body and how long will it take before you have 25 mg of caffenine in your body

OpenStudy (anonymous):

\[A = Pe^{rt}\] r = 15.6% or 0.156 \[50mg = 100mge^{rt} => 1/2=e^{rt}\] take the ln of both sides \[ln(1/2)/r = t= > ln(0.5)/0.156 = 4.443hours \]

OpenStudy (anonymous):

or \[1/2*P = Pe^{rt} \]

OpenStudy (anonymous):

For 25mg do \[1/4*P = Pe^{rt}\]

OpenStudy (toxicsugar22):

i dont know

OpenStudy (toxicsugar22):

can u show me the first time so i know hw the next time